Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.00 coming 3 years from today. The dividend should grow rapidly - at a rate of 33% per year - during Years 4 and 5; but after Year 5, growth should be a constant 5% per year.

If the required return on Computech is 15%, what is the value of the stock today? Round your answer to the nearest cent. Do not round your intermediate calculations.
Answer #1

The total present value of dividend till 5 years is calculated below:

1x 1.33 Present value = 1153 + 1.154T 1x1.334 1.155

Present value = 0.657516 +0.7604318 +0.879456

Present value = 2.2974

The present value of dividends receivable after 5 years can be calculated as:

Dividend x (1 + Growth rate) Required return – Growth rate ((1+ Required return)

Present value = 1.332 x (1 + 0.05) 0.15 -0.05 1.155

1.857345 Present value = = - X 0.4971767 0.10

Present value = 9.2343

THe stock value today is 9.2343+ 2.2974 = $11.53 (rounded off).